# Entanglement Entropy of Free Fermions in Timelike Slices

### v3

@@ IV Scaling Behavior and Oscillation Period at $\tau\leq\tau_{0}$
We now show the scaling behavior and oscillation period of the zero-temperature entanglement entropy $S_{A}$ in the time-direction slice with fixed time separation $\tau=\tau_{0},\frac{2}{\pi}\tau_{0}$ or $\frac{1}{2}\tau_{0}$. For $\tau\leq\tau_{0}$, the zero temperature $S_{A}$ is in general a linear function in $\ln K$ plus an oscillation in $K$. The oscillation amplitude is small when $\tau=\tau_{0}$ (Fig. S2), but is large when $\tau\ll\tau_{0}$ (Figs. S3 and S4). We first fit a linear function $S_{\text{fit}}=a\cdot\ln(K)+b$ to extract $a=\left(\frac{\partial S_{A}}{\partial\ln(K)}\right)_{\tau}$. Then we find the oscillation period of $S_{A}-S_{\text{fit}}$ (using the fitted parameters $a$ and $b$) against $K$. The oscillation is periodic in $K$. To mitigate the effect of oscillation to the fitting, we have excluded the entropy at small $K$, see captions of Tabs. 1, 2 and 3. The number of filled single-fermion eigenstates are $101,151,201,251,301,351$ for $\nu=0.2,0.3,0.4,0.5,0.6,0.7$ respectively (total size $L=500$), to avoid breaking the double degeneracy in the cases of cosine, $k^{2}$, and nonchiral spectrum in Eqs. S14, S15 and S16. The nonchiral spectrum in Eq. S16 gives almost the same entanglement entropy as chiral spectrum in Eq. S17 due to their nearly identical density of states $\Omega(E)$, so we have only plotted the entanglement entropy of the cosine, $k^{2}$ and chiral spectra in Figs. S2, S3 and S4.

### v4

@@ IV Scaling Behavior and Oscillation Period at $\tau\leq\tau_{0}$
We now show the scaling behavior and oscillation period of the zero-temperature entanglement entropy $S_{A}$ in the time-direction slice with fixed time separation $\tau=\tau_{0},\frac{2}{\pi}\tau_{0}$ or $\frac{1}{2}\tau_{0}$. For $\tau\leq\tau_{0}$, the zero temperature $S_{A}$ is in general a linear function in $\ln K$ plus an oscillation in $K$. The oscillation amplitude is small when $\tau=\tau_{0}$ (Fig. S2), but is large when $\tau\ll\tau_{0}$ (Figs. S3 and S4). We first fit a linear function $S_{\text{fit}}=a\cdot\ln(K)+b$ to extract $a=\left(\frac{\partial S_{A}}{\partial\ln(K)}\right)_{\tau}$. Then we find the oscillation period of $S_{A}-S_{\text{fit}}$ (using the fitted parameters $a$ and $b$) against $K$. The oscillation is periodic in $K$. To mitigate the effect of oscillation to the fitting, we have excluded the entropy at small $K$, see captions of Tabs. 1, 2 and 3. The number of filled single-fermion eigenstates are $101,151,201,251,301,351$ for $\nu=0.2,0.3,0.4,0.5,0.6,0.7$ respectively (total size $L=500$), to avoid breaking the double degeneracy in the cases of cosine, $k^{2}$, and nonchiral spectrum in Eqs. S33, S34 and S35. The nonchiral spectrum in Eq. S35 gives almost the same entanglement entropy as chiral spectrum in Eq. S36 due to their nearly identical density of states $\Omega(E)$, so we have only plotted the entanglement entropy of the cosine, $k^{2}$ and chiral spectra in Figs. S2, S3 and S4.